Title

INAF queue observing, 2014 Dec 19-20

Observer: F. Cusano, A. Marchetti
Telescope Operator: G. Bechetti
Support Astronomer: O. Kuhn
Instrument: MODS

Summary

Clear sky at sunset, humidity 10%.
The TO cannot open the dome at sunset because of snow on the roof.
The snow melted mostly on the left side of the roof, we can not open the right side
that is still covered with snow. We can use only LBT-left.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Saracco_B RDCS red grating 3.5h 1.5h through clouds
Standard gd71 red grating 9m  
4:22 The TO opens the dome (left part only)

5.00 problems with pointing (there is a problem with the guide star, which appears 3 mags fainter than it is) and mask selection (mask in the cassette positions 13, 14, 15 and 16 for low elevations, like in this case, is not selected). The mask problem is solved by pointing the telescope at zenith and manually selecting the mask

5:30 acqMODS gd71.acq, standard star
  • mods1r.20141220.0002-0004

5:39 execMODS gd71.obs
  • mods1r.20141220.0005-0007

5:41 doing again the darks for LUCI
  • luci.20141220.0002-0022
The last frame presents hot pixels blooming effect, so it's better to avoid using it for the reductions

5:56 acqMODS mods.7.RDCSm1 .acq, acquisition of Saracco_B target
  • still problems with guiding star and mask. We try to solve them with John Hill

6:03 acqMODS mods.7.RDCSm1.acq, acquisition of Saracco_B target
  • same magnitude problem as before with the guide star. The problem is in the probe filter wheel: if positioned on the clear filter, something is not working properly and the star gets filtered. We solve it by changing filter and using the F525LP. We want to see if homing the probe changes something

6:23 trying agin the acquisition for Saracco_B, acqMODS mods.7.RDCSm1.acq
  • we solved the filter problem, but the mask problem presented again. We have to move the telescope to zenith to manually select the mask

6:27 again the acquisition for Saracco_B, acqMODS mods.7.RDCSm1.acq. Problems: the mirror has been experimenting dramatic temperature changes, so there are collimation problems.
  • mods1r.20141220.0010-0011

7:30 we send the preset-only again for Saracco_B, acqMODS -p mods.7.RDCSm1.acq. Problems. We cancel that preset and send another one

7.34 Saracco_B, acqMODS -p mods.7.RDCSm1.acq. Still problems. The TO tries restarting TCS. Nothing is responding to whatever TO does on TCS

8:06 We send the acquisition again acqMODS mods.7.RDCSm1.acq, but recovering the offset from the previous onhe. Seeing is 0.8!
  • mods1r.20141220.0012-0014

8:23 We run the science of Saracco_B scientific target after correcting for some errors in the ob, execMODS mods.7.RDCSm1.obs
  • mods1r.20141220.0015-0018 clouds during the last half hour

9.50 We run again the acquisition for the second ob of Saracco_B, acqMODS mods.7.RDCSm1_2.acq , because, due to an error in the script, the offsets in the previous ob where strange
  • mods1r.20141220.0020-0022 clouds

9.55 We run the scientific integration, execMODS mods.7.RDCSm1_2.obs
  • mods1r.20141220.0023-0026 clouds

10.15 The TO has to unwrap the telescope

10.20 we recover the acquisition
  • mods1r.20141220.0027 clouds

10.35 The sky is looking better. We repeat execMODS mods.7.RDCSm1_2.obs
  • mods1r.20141220.0028-0031
Peaks of seeing 0.90!

12.45 It's twilight. We have time to run 2 exposures (30min total) on 4 of a new Saracco_B field: execMODS -e mods.7.RDCSm1_3.obs (we could have used the -e option also for the previous ob)
  • mods1r.20141220.0032-0033

1:25 the TO closes the dome

-- %USERSIG{FeliceCusano - 2014-12-19}%

Comments

 
Topic revision: r7 - 23 Jan 2015, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback